# Jina Embeddings 2: $8192$-Token General-Purpose Text Embeddings for Long Documents

### 02 Jan 2024 v2

@@ 1 Introduction
This paper is structured as follows: We begin with an overview of related work in Section 2. This is followed by an outline of the training process in Section 3, a description of the backbone model and its pre-training in Section 4, and a detailed walkthrough of the fine-tuning process for embeddings generation in Section 5. We culminate with an exhaustive evaluation in Section 6 and conclusions in Section 7.
@@ 2 Related Work
The E5 Wang et al. (2022), Jina  Embeddings v1 Günther et al. (2023), and GTE Li et al. (2023) collections of embedding models represent another leap forward. These models propose a holistic framework tailored for effective training across a myriad of tasks. This framework adopts a multi-stage contrastive training approach. An initial phase focuses on unsupervised training using a vast collection of weak pairs sourced from public data, enhancing the model’s domain generalization. Following this, a supervised fine-tuning stage employs a curated set of annotated text triples, representing diverse tasks. Together, these sequential stages yield state-of-the-art outcomes on the MTEB benchmark.
@@
3 Training Process Overview
The training process for Jina  Embeddings v2 is divided into three stages:
@@ 3 Training Process Overview
    Pre-training the Backbone: For the backbone pre-training, we design a modified BERT model capable of encoding documents with up to $8192$ tokens. This model is trained on a full-text corpus using a masked language modeling objective.
@@
    First Fine-tuning with Text Pairs: To encode a text passage into a single vector representation, the model is fine-tuned in an unsupervised manner on text pairs.
@@ 3. III
    Second Fine-tuning with Hard Negatives: The model is further fine-tuned using text pairs complemented with hard negatives. This stage is crucial for enabling the model to better distinguish between relevant passages and related, but irrelevant text passages.
@@
4 Backbone Pre-training
For the backbone training, we introduce a novel transformer model. Although its architecture is akin to the BERT model proposed by  Devlin et al. (2019), we implement several modifications to enhance its ability to encode extended text sequences and to generally bolster its language modeling capabilities. For the training process, we largely adopt the approach described in Liu et al. (2019a), incorporating additional performance optimizations.
@@ 4.1 Model Architecture
Model
Layers
Hidden
Params
jina-bert-small
4
512
33M
jina-bert-base
12
768
137M
jina-bert-large
24
1024
455M
@@
4.2 Full-Text Datasets for Pre-training
@@ 4.3 Training Algorithm
[…]
@@
Mirroring the optimization strategy of RoBERTa Liu et al. (2019a), we employ the AdamW algorithm Loshchilov and Hutter (2017), characterized by parameters $\beta_{1}=0.9$, $\beta_{2}=0.98$, $\epsilon=1e-6$, a weight decay of $0.01$, dropout set at $0.1$, and attention dropout also at $0.1$. Our learning rate schedule is linear, starting at 0 and peaking at a rate of $\eta$ post $10,000$ steps. Here, the values of $\eta$ are designated as $1\mathrm{e}{-3}$, $6\mathrm{e}{-4}$, and $4\mathrm{e}{-4}$ for the small, base, and large models respectively. A linear decay to zero ensues after reaching the $100,000$ steps threshold.
@@ 5 Fine-Tuning for Embeddings
After pre-training the Jina BERT models, we further fine-tune each of the models to encode a text sequence into a single vector representation. The core idea behind our embedding approach is inspired by the Sentence-BERT technique Reimers and Gurevych (2019). To enable a model to perform a text operation, we augment it with a mean pooling layer. This mean pooling step averages the token embeddings to merge their information into a single representation, without introducing additional trainable parameters. The training process for this enhanced model consists of an unsupervised phase followed by a supervised one.
5.1 First Unsupervised Fine-tuning
[…]
Construction of the Text-Pair Datasets:
@@ 5.1 First Unsupervised Fine-tuning
Training:
The goal of unsupervised fine-tuning is to encode text values that constitute a pair into analogous embedding representations, while encoding texts that aren’t paired into distinct embeddings. To achieve this contrastive goal, we employ the InfoNCE (van den Oord et al., 2018) loss function, similar to our earlier embedding models Günther et al. (2023). This loss function calculates the loss value for a pair $(q,p)\sim B$ within a batch $B\in D^{k}$ of $k$ text pairs as follows:
$\mathcal{L}_{\mathrm{NCE}}^{\mathrm{pairs}}(B):=\mathbb{E}_{(q,p)\sim B}\left[-\ln\frac{e^{s(q,p)/\tau}}{\sum\limits_{i=1}^{k}e^{s(q,p_{i})/\tau}}\right]$ (3)
@@
$\mathcal{L}^{\mathrm{pairs}}(B)$ $:=\mathcal{L}^{\mathrm{pairs}}_{\mathrm{NCE}}(B)+\mathcal{L}^{\mathrm{pairs}}_{\overline{\mathrm{NCE}}}(B),\text{ with}$
$\mathcal{L}_{\overline{\mathrm{NCE}}}^{\mathrm{pairs}}(B)$ $:=\mathbb{E}_{(q,p)\sim B}\left[-\ln\frac{e^{s(p,q)/\tau}}{\sum\limits_{i=1}^{k}e^{s(p,q_{i})/\tau}}\right]$ (4)
@@
5.2 Second Supervised Fine-tuning
@@
Dataset with annotated negatives:
@@ 5.3 Memory Optimizations
When training embedding models, having a large batch size is crucial. This is because the InfoNCE loss functions $\mathcal{L}^{pairs}$ and ${L}_{\mathrm{NCE}^{+}}$ compute the loss values based on the entirety of the batch. The batch size determines the number of text values each individual text value is compared against. As a result, the computed loss value might not be as expressive with smaller batches. Li et al. (2023) provided an in-depth analysis, highlighting the positive impact of larger batch sizes on the performance of the resultant embedding model. To accommodate larger batch sizes, it becomes essential to minimize the memory overhead during training. We achieved this by training our models in mixed precision Micikevicius et al. (2018) and leveraging the deepspeed Rasley et al. (2020) framework for further optimization. Activation checkpointing Chen et al. (2016) was also employed to curtail memory usage. Specifically, we inserted a checkpoint after each BERT layer within our model.
@@ 6.1 Evaluation of Jina BERT
Figure 2: Variation of model MLM accuracy in function of the sequence length
@@ 6.2.1 Results on MTEB
text-embedding-ada-002 unknown 70.93 45.9 84.89 56.32 49.25 80.97 30.8 60.99
e5-base-v2 110M 73.84 43.8 85.73 55.91 50.29 81.05 30.28 61.5
all-MiniLM-L6-v2 23M 63.05 42.35 82.37 58.04 41.95 78.9 30.81 56.26
@@
jina-small-v2 33M 68.82 40.08 84.44 55.09 45.14 80 30.56 58
jina-base-v2 137M 73.45 41.73 85.38 56.98 47.87 80.7 31.6 60.38
@@
Figure 3: Evaluation with Different Maximum Sequence Lengths [^9] For e5-base-v2, we abstained from employing specific prefixes like “query: ”, which might result in varied evaluation outcomes. text-embedding-ada-002 caps its encoding at 8191 tokens, not $8192$.
@@ 6.2.2 Impact of Maximum Sequence Length
To discern how sequence length impacts the accuracy of downstream tasks, we executed long document clustering and retrieval tasks, modulating the tokenizer’s maximum sequence length. This allows us to gauge the models’ performance on variable sequence lengths through truncation. Since a majority of the extant tasks in the MTEB feature documents under $512$ tokens, we resort to our three novel datasets elucidated in Section 6.2, accessible on Hugging Face [^10] Our datasets are available at https://huggingface.co/jinaai. Furthermore, we employ the SciFact dataset Wadden et al. (2020), given its substantial count of texts exceeding $512$ tokens.
@@ 7 Conclusion
Appendix A Appendix
@@
Task
jina-small-v2 jina-base-v2
AmazonCounterfactualClassification
71.36 74.73
AmazonPolarityClassification
82.9 88.54
AmazonReviewsClassification
40.89 45.26
Banking77Classification
78.25 84.01
EmotionClassification
44.01 48.77
ImdbClassification
73.64 79.44
MassiveIntentClassification
67.61 71.93
MassiveScenarioClassification
69.75 74.49
MTOPDomainClassification
93.96 95.68
MTOPIntentClassification
72.50 83.15
ToxicConversationsClassification
71.54 73.35
TweetSentimentExtractionClassification
59.40 62.06
Avg
68.82 73.45
@@
Task
jina-small-v2 jina-base-v2
ArxivClusteringP2P
44.02 45.39
ArxivClusteringS2S
35.16 36.68
BiorxivClusteringP2P
35.57 37.05
BiorxivClusteringS2S
29.07 30.16
MedrxivClusteringP2P
31.86 32.41
MedrxivClusteringS2S
27.51 28.09
RedditClustering
49.28 53.05
RedditClusteringP2P
57.09 60.31
StackExchangeClustering
55.35 58.52
StackExchangeClusteringP2P
34.42 34.96
TwentyNewsgroupsClustering
41.57 42.47
Avg
40.08 41.73
@@
Task
jina-small-v2 jina-base-v2
SummEval
30.56 31.6
@@
Task
jina-small-v2 jina-base-v2
SprintDuplicateQuestions
95.12 95.30
TwitterSemEval2015
72.15 74.74
TwitterURLCorpus
86.05 86.09
Avg
84.44 85.38
@@
Task
jina-small-v2 jina-base-v2
AskUbuntuDupQuestions
59.62 62.25
MindSmallReranking
30.99 30.54
SciDocsRR
79.76 83.10
StackOverflowDupQuestions
49.99 52.05
Avg
55.09 56.98
@@
Task
jina-small-v2 jina-base-v2
ArguAna
46.73 44.18
ClimateFEVER
20.05 23.53
CQADupstackRetrieval
38.03 39.34
DBPedia
32.65 35.05
FEVER
68.02 72.33
FiQA2018
33.43 41.58
HotpotQA
56.48 61.38
MSMARCO
37.28 40.92
NFCorpus
30.40 32.45
NQ
51.59 60.44
QuoraRetrieval
87.19 88.20
SCIDOCS
18.61 19.86
SciFact
63.89 66.68
Touche2020
23.52 26.24
TRECCOVID
65.18 65.91
Avg
45.14 47.87
@@
Task
jina-small-v2 jina-base-v2
BIOSSES
80.52 81.23
SICK-R
76.72 79.65
STS12
73.66 74.27
STS13
83.30 84.18
STS14
79.17 78.81
STS15
87.30 87.55
STS16
83.61 85.35
STS17(en-en)
88.23 88.88
STS22(en)
63.46 62.20
STSBenchmark
84.04 84.84
Avg
80.00 80.70

### 04 Feb 2024 v4

@@ 1 Introduction
This paper is structured as follows: We begin with an overview of related work in Section 2. This is followed by an outline of the training paradigm in Section 3, a description of the backbone model and its pre-training in Section 4, and a detailed walkthrough of the fine-tuning process for embeddings generation in Section 5. We culminate with an exhaustive evaluation in Section 6 and conclusions in Section 7.
@@ 2 Related Work
The E5 Wang et al. (2022), Jina  Embeddings v1 Günther et al. (2023), and GTE Li et al. (2023) collections of embedding models represent another leap forward. These models propose a holistic framework tailored for effective training across a myriad of tasks. This framework adopts a multi-stage contrastive training approach. An initial phase focuses on training using a vast collection of weak pairs sourced from public data, enhancing the model’s domain generalization. Following this, a supervised fine-tuning stage employs a curated set of annotated text triples, representing diverse tasks. Together, these sequential stages yield state-of-the-art outcomes on the MTEB benchmark.
@@
3 Training Paradigm Overview
The training paradigm for Jina  Embeddings v2 is divided into three stages:
@@ 3 Training Paradigm Overview
    Pre-training a Modified BERT: For the backbone language model, we propose a modified BERT model capable of encoding documents with up to $8192$ tokens. This model is trained from scratch on a full-text corpus using a masked language modeling objective.
@@
    Fine-tuning with Text Pairs: To encode a text passage into a single vector representation, the model is fine-tuned on text pairs.
@@ 3. III
    Fine-tuning with Hard Negatives: The model is further fine-tuned using text pairs complemented with hard negatives. This stage is crucial for enabling the model to better distinguish between relevant passages and related, but irrelevant text passages.
@@
4 Pre-training a Modified BERT
For the backbone language model, we introduce a novel transformer based on BERT Devlin et al. (2019) with several modifications to enhance its ability to encode extended text sequences and to generally bolster its language modeling capabilities. For the training process, we largely adopt the approach described in Liu et al. (2019a), incorporating additional performance optimizations.
@@ 4.1 Model Architecture
Model Layers Hidden Params
Jina BERT Small 4 512 33M
Jina BERT Base 12 768 137M
Jina BERT Large 24 1024 455M
@@
4.2 Training Data
@@ 4.3 Training Algorithm
The masked tokens are predicted by a decoder $f:\mathbb{R}^{d}\to\mathbb{R}^{|V|}$, which takes the output token embedding $\bm{e_{i}}\in\mathbb{R}^{d}$ of a masked token and predicts a probability for each token in the vocabulary. The loss $L_{\mathrm{MLM}}$ is computed by evaluating the cross entropy between the predicted probabilities and the actual masked tokens, as described in Equation (2). Here, $I:\{1,\ldots,n\}\to|V|$ denotes the function that maps each of the $n$ masked tokens to its respective index in the vocabulary:
$\mathcal{L}_{\mathrm{MLM}}(t)$ $:=\sum\limits_{k=1}^{n}\ln f(\bm{e_{i}})_{I(k)}$ (2)
@@
Mirroring the optimization strategy of RoBERTa Liu et al. (2019a), we employ the AdamW algorithm Loshchilov and Hutter (2017), characterized by parameters $\beta_{1}=0.9$, $\beta_{2}=0.98$, $\epsilon=1\mathrm{e}{-6}$, a weight decay of $0.01$, dropout set at $0.1$, and attention dropout also at $0.1$. Our learning rate schedule is linear, starting at 00 and peaking at a rate of $\eta$ post $10,000$ steps. Here, the values of $\eta$ are designated as $1\mathrm{e}{-3}$, $6\mathrm{e}{-4}$, and $4\mathrm{e}{-4}$ for the small, base, and large models respectively. A linear decay to zero ensues after reaching the $100,000$ steps threshold.
@@ 5 Fine-Tuning for Embeddings
After pre-training the Jina BERT models, we further fine-tune each of the models to encode a text sequence into a single vector representation. The core idea behind our embedding approach is inspired by the Sentence-BERT Reimers and Gurevych (2019). To enable a model to perform a text operation, we augment it with a mean pooling layer. This mean pooling step averages the token embeddings to merge their information into a single representation, without introducing additional trainable parameters. The training process for this enhanced model consists of an unsupervised phase followed by a supervised one.
5.1 Fine-tuning with Text Pairs
During the first fine-tuning stage, we train the models on a corpus of text pairs $(q,p)\in\mathbb{D}^{\mathrm{pairs}}$, comprising a query string $q$ and a target string $p$.
Training Data
@@ 5.1 Fine-tuning with Text Pairs
Loss Function:
The goal of this fine-tuning stage is to encode text values that constitute a pair into analogous embedding representations, while encoding texts that aren’t paired into distinct embeddings. To achieve this contrastive goal, we employ the InfoNCE (van den Oord et al., 2018) loss function, similar to our earlier embedding models Günther et al. (2023). This loss function calculates the loss value for a pair $(q,p)\sim\mathbf{B}$ within a batch $\mathbf{B}\subset\mathbb{D}^{\mathrm{pairs}}$ as follows:
$\mathcal{L}_{\mathrm{NCE}}^{\mathrm{pairs}}(\mathbf{B}):=\mathbb{E}_{(q,p)\sim\mathbf{B}}\left[-\ln\frac{e^{s(q,p)/\tau}}{\sum\limits_{i=1}^{k}e^{s(q,p_{i})/\tau}}\right]$ (3)
@@
$\mathcal{L}^{\mathrm{pairs}}(\mathbf{B})$ $:=\mathcal{L}^{\mathrm{pairs}}_{\mathrm{NCE}}(\mathbf{B})+\mathcal{L}^{\mathrm{pairs}}_{\overline{\mathrm{NCE}}}(\mathbf{B}),\text{ with}$
$\mathcal{L}_{\overline{\mathrm{NCE}}}^{\mathrm{pairs}}(\mathbf{B})$ $:=\mathbb{E}_{(q,p)\sim\mathbf{B}}\left[-\ln\frac{e^{s(p,q)/\tau}}{\sum\limits_{i=1}^{k}e^{s(p,q_{i})/\tau}}\right]$ (4)
@@
5.2 Fine-tuning with Hard Negatives
@@
Training Data
@@ 5.3 Memory Optimizations
When training embedding models, having a large batch size is crucial. This is because the InfoNCE loss functions $\mathcal{L}^{\mathrm{pairs}}$ and ${L}_{\mathrm{NCE}^{+}}$ compute the loss values based on the entirety of the batch. The batch size determines the number of text values each individual text value is compared against. As a result, the computed loss value might not be as expressive with smaller batches. Li et al. (2023) provided an in-depth analysis, highlighting the positive impact of larger batch sizes on the performance of the resultant embedding model. To accommodate larger batch sizes, it becomes essential to minimize the memory overhead during training. We achieved this by training our models in mixed precision Micikevicius et al. (2018) and leveraging the deepspeed Rasley et al. (2020) framework for further optimization. Activation checkpointing Chen et al. (2016) was also employed to curtail memory usage. Specifically, we inserted a checkpoint after each BERT layer within our model.
@@ 6.1 Evaluation of Jina BERT
Figure 2: Variation of model MLM accuracy w.r.t. the sequence length
@@ 6.2 Evaluation of Jina  Embeddings v2
We further evaluated Jina  Embeddings v2 using a novel benchmark, referred to as LoCo  [^9] https://hazyresearch.stanford.edu/blog/2024-01-11-m2-bert-retrieval. The LoCo dataset consists of five retrieval tasks derived from publicly available datasets. The selection process for these tasks was guided by several criteria, notably the length of the documents, with a preference towards longer texts, in addition to a manual review to verify that the tasks require a thorough understanding of the entire document. The results of our models on the LoCo dataset are provided in Table 11.
@@ 6.2.1 Results on MTEB
text-embedding-ada-002 unknown 70.93 45.90 84.89 56.32 49.25 80.97 30.80 60.99
e5-base-v2 110M 73.84 43.80 85.73 55.91 50.29 81.05 30.28 61.50
all-MiniLM-L6-v2 23M 63.05 42.35 82.37 58.04 41.95 78.90 30.81 56.26
@@
jina-small-v2 33M 68.82 40.08 84.44 55.09 45.64 80.00 30.56 58.12
jina-base-v2 137M 73.45 41.74 85.38 56.99 48.45 80.70 31.60 60.37
@@
Figure 3: Evaluation w.r.t. maximum sequence length. For e5-base-v2, we abstained from employing specific prefixes like query: , which might result in varied evaluation outcomes. Note, text-embedding-ada-002 caps its context length at $8191$ tokens, not $8192$.
@@ 6.2.2 Impact of Maximum Sequence Length
To discern how sequence length impacts the accuracy of downstream tasks, we executed long document clustering and retrieval tasks, modulating the tokenizer’s maximum sequence length. This allows us to gauge the models’ performance on variable sequence lengths through truncation. Since a majority of the extant tasks in the MTEB feature documents under $512$ tokens, we resort to our three novel datasets elucidated in Section 6.2, accessible on Hugging Face. Furthermore, we employ the SciFact dataset Wadden et al. (2020), given its substantial count of texts exceeding $512$ tokens.
@@ 7 Conclusion
Appendix A Appendix: MTEB and LoCo Becnharmk
@@
Task jina-small-v2 jina-base-v2
AmazonCounterfactualClassification 71.36 74.73
AmazonPolarityClassification 82.90 88.54
AmazonReviewsClassification 40.89 45.26
Banking77Classification 78.25 84.01
EmotionClassification 44.01 48.77
ImdbClassification 73.64 79.44
MassiveIntentClassification 67.61 71.93
MassiveScenarioClassification 69.75 74.49
MTOPDomainClassification 93.96 95.68
MTOPIntentClassification 72.50 83.15
ToxicConversationsClassification 71.54 73.35
TweetSentimentExtractionClassification 59.40 62.06
Avg 68.82 73.45
@@
Task jina-small-v2 jina-base-v2
ArxivClusteringP2P 44.02 45.39
ArxivClusteringS2S 35.16 36.68
BiorxivClusteringP2P 35.57 37.05
BiorxivClusteringS2S 29.07 30.16
MedrxivClusteringP2P 31.86 32.41
MedrxivClusteringS2S 27.51 28.09
RedditClustering 49.28 53.05
RedditClusteringP2P 57.09 60.31
StackExchangeClustering 55.35 58.52
StackExchangeClusteringP2P 34.42 34.96
TwentyNewsgroupsClustering 41.57 42.47
Avg 40.08 41.73
@@
Task jina-small-v2 jina-base-v2
SummEval 30.56 31.60
@@
Task jina-small-v2 jina-base-v2
SprintDuplicateQuestions 95.12 95.30
TwitterSemEval2015 72.15 74.74
TwitterURLCorpus 86.05 86.09
Avg 84.44 85.38
@@
Task jina-small-v2 jina-base-v2
AskUbuntuDupQuestions 59.62 62.25
MindSmallReranking 30.99 30.54
SciDocsRR 79.76 83.10
StackOverflowDupQuestions 49.99 52.05
Avg 55.09 56.98
@@
Task jina-small-v2 jina-base-v2
ArguAna 46.73 44.18
ClimateFEVER 20.05 23.53
CQADupstackRetrieval 38.03 39.34
DBPedia 32.65 35.05
FEVER 68.02 72.33
FiQA2018 33.43 41.58
HotpotQA 56.48 61.38
MSMARCO 37.28 40.92
NFCorpus 30.40 32.45
NQ 51.59 60.44
QuoraRetrieval 87.19 88.20
SCIDOCS 18.61 19.86
SciFact 63.89 66.68
Touche2020 23.52 26.24
TRECCOVID 65.18 65.91
Avg 45.14 47.87
@@
Task jina-small-v2 jina-base-v2
BIOSSES 80.52 81.23
SICK-R 76.72 79.65
STS12 73.66 74.27
STS13 83.30 84.18
STS14 79.17 78.81
STS15 87.30 87.55
STS16 83.61 85.35
STS17(en-en) 88.23 88.88
STS22(en) 63.46 62.20
STSBenchmark 84.04 84.84
Avg 80.00 80.70
@@
Model Fine-tuned on LoCo Parameters Context Length avg. nDCG@10
M2-BERT-32768 80M 32,768 92.5
e5-mistral-7b-instruct 7.3B 4,096 88.5
M2-BERT-32768 80M 8,192 85.9
jina-base-v2 137M 8192 85.4
bge-large-en-v1.5 335M 512 85.0
M2-BERT-2048 80M 2,048 83.6
jina-small-v2 33M 8,192 83.4
bge-base-en-v1.5 109M 512 83.0
bge-small-en-v1.5 33M 512 81.2
bge-large-en-v1.5 335M 512 77.2
bge-base-en-v1.5 109M 512 73.4
bge-small-en-v1.5 33M 512 70.6
cohere-embed-v3 NA 512 66.6
ada-embeddings-002 NA 8,191 52.7
voyage-v1 NA 4,096 25.4
Table 11: Performance on the new LoCo Dataset